OpenStudy (anonymous):

Sharon earns $25 per item she sells plus a base salary of $100 per week. Write and solve an inequality to find how many items she must sell to earn at least $700 per week. Provide your conclusion as a complete sentence.

OpenStudy (goldphenoix):

100+25x= 700

OpenStudy (anonymous):

(Let i represent the number of items sold) \[$25p +$100 \ge $700\] now subtract 100 from both sides, now you are left with \[$25p \ge $600\], divide both sides by 25 to get p (the least amount of items Sharon sells) \[\frac{ $25p }{ 25 } \ge \frac{ $600 }{ 25 }\] If I am wrong correct me.

OpenStudy (goldphenoix):

Think about it. She constantly earns 100 dollars per week. But she also earn 25 dollars per things she sell. The things she sell is not constant. So you can put a variable. 100 dollars + 25x = 700; where x is the amount of things she sold. Do you understand? I hope this help!

OpenStudy (goldphenoix):

Oh yeah, and solve the inequality, 100 + 25x Greater than or equal to 700

OpenStudy (goldphenoix):

Eventually after you did the inequality sucessfully, you should get x Greater than or equal to 24. So 24,25,26,27,100, etc. should give her 700 dollars or more per week.

OpenStudy (texaschic101):

@justin12lo is correct....and so is @GoldPhenoix ....600/25 = x....now just divide to find x.

OpenStudy (anonymous):

600/25=575

OpenStudy (anonymous):

Wait did you subtract 600 and 25 to get 575, because the / means divide 600 and 25 to get x is greater than or equal to 24. When you plug it in, you get 25(24) + 100 + 700, which is 600 + 100 = 700.

OpenStudy (anonymous):

I mean greater than or equal to